OpenStudy (anonymous):

pleasee simplify!!! √-64/(7-6i)-(2-2i)

OpenStudy (lgbasallote):

is that \[\large \frac{\sqrt{-64}}{7 - 6i} - 2 - 2i?\]

OpenStudy (anonymous):

/((7-6i)-(2-2i))

OpenStudy (lgbasallote):

oh so \[\frac{\sqrt{-64}}{7-6i-(2-2i)}??\]

OpenStudy (anonymous):

(7-6i)-(2-2i) is the denominator and the nominator is the square root of -64